OpenStudy (anonymous):

Write each percent as a fraction or a mixed number in simplest form. 36%, 33 1/3%, 124%, 27%.

OpenStudy (anonymous):

please help me

OpenStudy (gorv):

36%=36/100=9/25

OpenStudy (anonymous):

ok

OpenStudy (gorv):

\[33\frac{ 1 }{ 3 }%=100/(3*100)=1/3\]

OpenStudy (gorv):

124/100=31/25

OpenStudy (gorv):

27/100

OpenStudy (anonymous):

ok i dont think that you can simplify that